# Evaluation of Language Models in the Medical Context Under Resource-Constrained Settings

### v1

@@ 5 Results and Discussion
The main findings are outlined below, each supported by the figures and tables provided. These findings can be further verified by the figures in Appendix C.
5.1 Model size: more parameters do not generally translate into better performance
The experiments conducted for text classification yield findings that challenge the assertion that larger models consistently deliver superior performance. An analysis of the impact of the logarithm of size on model performance indicates insufficient evidence to conclude that the Pearson correlation between size and performance is statistically significant in all the cases (refer to Fig. 2). This observed discrepancy may be attributed to several factors, including the relatively small size of the models considered, with a number of parameters up to billions rather than tens or hundreds of billions.
Specifically, in contextual embedding similarity, the trend of performance improvement with increasing size is almost nonexistent. As depicted in Fig. 3, for instance, models such as SapBERT (m06) and BioLORD (m07-m09), which excel in this approach, outperform models even three times larger in both datasets. When solely the performance deltas associated with increasing the number of parameters for the same type of model are examined, they are both positive and negative. For the transcriptions dataset, as indicated by the lighter crosses in Fig. 3, BERT (m01) marginally outperforms BERT (m00), similar to the case of BiomedBERT-large (abstracts only) (m04) versus BioMedBERT (abstracts only) (m03) where positive deltas are also obtained. However, for the MS-CXR dataset, BERT (m00) surpasses BERT (m01) in all metrics, and BioMedBERT (abstracts only) (m03) outperforms BiomedBERT-large (abstracts only) (m04) in precision. Now, by comparing BioMedBERT (abstracts only) (m03) and BiomedBERT (abstracts + full text) (m02), performance gains are obtained when more training data are used.
In natural language inference, while the results align with the expectation that larger models tend to lead to better performance, more experiments are needed to draw a solid conclusion. In the MS-CXR dataset, the difference between the most effective NLI-DeBERTa model and the least effective RoBERTa-MNLI and BART Large-MNLI models ranges from across the metrics considered. These differences suggest that larger models among the models used always lead to performance improvement. However, reaching the same conclusion for the transcriptions dataset is not straightforward, given the results especially obtained for RoBERTa-MNLI (m12), which, as depicted in Fig. 3, is closer to the performance of NLI-DeBERTa (m11) than to that of BART Large-MNLI (m13).
When it comes to multiple-choice question-answering, the impact of the increase in size on performance is also unclear. While positive Pearson correlations are obtained, there is insufficient evidence to deem them statistically significant. When analyzing this impact within the T5 model family, as shown in Fig. 3, the effect is minimal or inconsistent when considering their non-instruction-tuned versions (m14-m17, m24-m25). Conversely, when examining the instruction-tuned versions of T5 (m18-m23), a consistent positive effect of size on performance is observed for the transcriptions dataset. Although a positive effect is also present for the MS-CXR dataset, it is only for T0 models (m22-m23).
Overall, the results of the text classification task suggest that increasing model size only sometimes leads to improved performance, whether comparing different models or evaluating the same models. Therefore, while model size may be a relevant factor in determining performance, training data and objectives may be more decisive. Expanding the sample size could be essential to validate this observation, considering a minimum of 30 or 35 models per approach. The previous insights can be verified in more detail in Figs. 13, 14, 15, 16, 17 and 18 in Appendix C.
In the context of the conditional text generation task, as shown in Fig. 4, increasing the size of the model contributes to higher performances, regardless of whether the BOS token is included. Also, an improvement in the performance of the second versions (m42, m44) versus the first versions (m41, m43) of OpenLLaMA is observed. Considering that the difference between these OpenLLaMA versions is the dataset used for pre-training, the conditional text generation task results do not contradict the above conclusion for text classification. Furthermore, our findings do not conflict with existing scaling laws, usually associated with LLMs. First, because we have no control over training, i.e., we cannot guarantee that the model is not exposed to variations other than the number of parameters (size). Second, Chinchilla’s law does not consider inter-model comparisons.
5.2 Model domain: more than a specialized domain; model architecture, training data, and training objective
Current medical datasets remain relatively small compared to those of the general domain, covering only a tiny region of the medical knowledge space [79]. Since some of the available domain-specific models are specialized by continuous pre-training on only one of these datasets, their generalization ability may be hampered, counter to the goal of creating well-generalizable domain-specific models and even achieving the ideal of artificial general intelligence (AGI). As discussed previously, the results suggest that architecture, training data, and training objectives are crucial in determining model generalization, possibly outweighing model size.
For the classification task through contextual embedding similarity, the effectiveness of domain specialization in improving performance is not evident, as displayed in Fig. 3. For the transcriptions dataset, refer to the lighter blue crosses in the figure. For example, Bio+Clinical BERT (m10) achieves lower results than expected, being around the middle of the performance ranking. Even in biomedical domain specialization, BERT (m00), a general domain model, outperforms some of the BiomedBERT models (m02-m04), challenging the superiority of domain-specific models. These observations are valid for both datasets. In particular, the latter observation also extends to BERT (m01).
Two comparisons arise when employing multiple-choice question-answering: T5 models (m14-m15) versus their clinical specialized versions (m24-m25) and Alpaca (m39) versus MedAlpaca (m52). However, the evidence supporting the hypothesis that domain specialization significantly improves performance is limited, as displayed in Fig. 3. Specifically, in the transcriptions dataset, differences in AUC scores between ClinicalT5 (m24-m25) and T5 (m14-m15) are and for the base and large sizes, respectively. This trend persists across all prompts. Similarly, the difference in AUC scores between MedAlpaca (m52) and Alpaca (m39) is , negatively affecting performance. In the MS-CXR dataset, the difference in F1-scores between ClinicalT5 (m24-m25) and T5 (m14-m15) is and for the base and large sizes, respectively. For MedAlpaca (m52) versus Alpaca (m39), the F1-score deteriorates even by . Thus, the evidence suggests that domain specialization provides either minimal or nonexistent benefits for this task.
Similar patterns to those discussed previously are identified in the text generation task, as observed in Fig. 4. Remarkably, the second version of LLaMA, LLaMA 2-7B (m38), stands out as the model with the highest predictive capacity for the dataset, with an average perplexity of . In contrast, BioGPT (m48) presents the greatest difficulty in comprehending the dataset, with an average perplexity of . These results are paradoxical, especially considering that BioGPT (m48) is domain-specialized while LLaMA 2-7B (m38) is not. A more detailed analysis by size reveals that
1.
for the medium size, BioGPT (m48) does not outperform any of the general domain models, whereas GPT-2-PubMed Medium (m46) does;
2.
for the large size, domain specialization proves beneficial; and
3.
for the XL and XXL sizes, neither Galactica (m50-m51) nor BioGPT-Large (m49) clearly outperforms general domain models.
Consequently, the only specialized models that demonstrate an advantage are GPT-2-PubMed (m46-m47).
Considering the insights discussed and the remarkable performance of SapBERT (m06) and BioLORD (m07-m09) models in the contextual embedding similarity approach, as well as Flan-T5 (m18-m21) and T0 (m22-m23) in the multiple-choice question-answering approach, the fundamental role played by training data, training objectives, and model architectures is underscored. Continual pre-training for named entity recognition or medical entity linkage using contrastive learning on UMLS data might contribute to the success of SapBERT and BioLORD models in the contextual embedding similarity task. On the other hand, employing instruction-tuned text-to-text models represents a compelling approach to achieving high performance in multiple-choice question-answering. Additionally, exploring not only the hypothesis but also the impact of domain instruction fine-tuning could provide further insights into model performance and effectiveness in specialized domains. Given that it is impossible to conclude the impact of the model domain on NLI, expanding the analysis to incorporate domain-specialized NLI models in biomedical and clinical domains could be valuable.
5.3 Prompting and instruction-tuned models: key in determining model performance
One of the central points of the study is to analyze the influence of prompts on the models and text classification approaches under investigation. As shown in Fig. 5, the results underscore the fundamental role of the prompt and its wording in the performance of the model, demonstrating both positive and negative impacts. Consequently, we advocate using prompts and advanced prompt engineering techniques to guide the model toward better results. Nevertheless, this process should not be limited to a single prompt due to the observed and well-known phenomenon of prompt brittleness [11]. As well, instruction fine-tuning proves advantageous for the models. Detailed results for each classification approach are presented below.
In contextual embedding similarity, using a prompt does not always confer benefits. For the transcripts dataset, the mean impact on the AUC score is points, with differences ranging from to points. Overall, using any of the proposed prompts improves performance for of the model + pooling strategy combinations. As for the BioLORD-PMB, BiomedBERT family, BERT, and SciBERT models, none of the suggested prompts led to AUC score improvements. Inversely, for the MS-CXR dataset, the prompt impact on performance is generally more positive, albeit with more variability. The mean impact, along with its range, is (, ) for accuracy, (, ) for F1-score, (, ) for precision, and (, ) for recall. Overall, employing any of the proposed prompts represents benefits for to of the model + pooling strategy combinations, depending on the metric. As for the BioMedBERT (abstracts only) and BiomedBERT-large (abstracts only), their performance scores, except for precision, improve with any of the proposed prompts, while for the BioLORD family and Bio+Clinical BERT, their scores are always hindered.
When examining the prompt impact on NLI, more consistent benefits are observed than for contextual embedding similarity. Using any of the proposed prompts yields improvements in the transcriptions dataset. The mean impact is for BART Large-MNLI, for NLI-DeBERTa, and for RoBERTa-MNLI, with larger models profiting the most from its usage. On its side, in the MS-CXR dataset, using a prompt only sometimes results in gains, particularly for NLI-DeBERTa. For this model, the impact on the F1-score is , while for BART Large-MNLI and RoBERTa-MNLI, it is and , respectively. For NLI-DeBERTa, positive prompt impacts are only observed for precision. For NLI, there are certain prompts with a high positive impact, whereas others do not, mostly independent of the model.
Regarding multiple-choice question-answering, the importance of prompts is evident, as they can significantly influence model performance. For example, in the transcriptions dataset, the proportion of models performing better than random guessing (AUC scores greater than ) can improve from to with appropriate prompts. Similarly, in the MS-CXR dataset, this proportion can change from to when measured by the F1-score (scores greater than ). Thus, its importance is highlighted not only by the high performance obtained but also by its variability, as reflected in Fig. 5. This is further supported by Figs. 17 and 18 in Appendix C. When analyzing the impact of templates versus questions, both components of the prompt (see Appendix B), it is found that the wording of templates has more influence on performance than the wording of questions, and no single prompt (template + question) works universally well for all models.
More specifically, in the transcriptions dataset, the performance ranking between the T5 models and between the Flan-T5 and T0 models remains consistent independent of the prompt. Moreover, Flan-T5-XXL (m21) and T0++ (m23) consistently outperform the other models regardless of the prompt. On average, the impact of the template, measured by the standard deviation of the AUC score per model, is , while the effect of the question within a template is . For the MS-CXR dataset, it becomes evident that some models, such as T5 and Falcon-7B, are unsuitable for this approach and dataset. The prompt plays an even more critical role than in the transcriptions dataset, affecting the ranking of models. The sensitivity is such that, with certain prompts, instruction-tuned models yield similar results to their base counterparts. On average, the impact of the template is , while the effect of the question within a template is . Interestingly, one of the proposed role-based prompts works remarkably well for the instruction-tuned versions of the T5 family despite not being the optimal choice overall.
Concerning instruction-tuning, these models generally outperform their non instruction-tuned counterparts. The instruction-tuned T5 versions, whether T0 or Flan-T5, in any size considered, exhibit superior performance than their base counterparts. In the transcriptions dataset, Flan-T5 either performs better or is on par with T0. However, in the MS-CXR dataset, T0 outperforms Flan-T5 in the 11B parameter version only when comparing the highest performances due to the considerable variability of T0++. Also, instruction-tuning improves performance consistently for the LLaMA group, whereas this is not always the case for the other generative models: MPT and GPT-J are exceptions in the transcriptions dataset and Falcon in the MS-CXR dataset. Overall, instruction-tuning represents a gain, with mean improvements of in AUC-score and in F1 score between the highest performances per model and in AUC-score and in F1 score between the mean performances for the transcriptions dataset and the MS-CXR dataset, respectively.
More details on the prompt impact can be found in Figs. 20 and 20 in Appendix C.
5.4 Text classification: instruction-tuned T5: models to go for
| CES | NLI | MCQA | ||||||||
| Dataset | Metric | Model (poolingprompt) | Score | Model (prompt) | Score | Model (prompt) | Score | |||
| Accuracy | BioLORD-STAMB2-v1-STS2 | (avg5) | 69.65 | RoBERTa-MNLI | (5) | 76.50 | T0++ | (t03+q04, t03+q05) | 81.68 | |
| F1-score | BioLORD-STAMB2-v1-STS2 | (avg ) | 69.34 | RoBERTa-MNLI | (5) | 78.25 | T0++ | (t03+q05) | 83.90 | |
| MS-CXR | Precision | BioLORD-PMB | (CLS ) | 83.34 | RoBERTa-MNLI | (5) | 80.74 | Alpaca 7B | (t02+q03) | 85.85 |
| Recall | BioLORD-STAMB2-v1-STS2 | (avg ) | 72.58 | RoBERTa-MNLI | (5) | 82.28 | T0++ | (t03+q04) | 89.18 | |
@@ 5.4 Text classification: instruction-tuned T5: models to go for
| Transcriptions | AUC score | BioLORD-STAMB2-v1-STS2 | (avg1) | 89.01 | BART Large-MNLI | (4) | 80.75 | Flan-T5-XXL | (t01+q09) | 92.38 |
@@
As shown in Table 2 and Fig. 3, for both datasets, the most favorable results are obtained with larger instruction-tuned T5 models, namely Flan-T5 (m18-m21) and T0 (m22-m23). The latter highlights the possibility of representing discriminative tasks as generative ones by framing them as instructions. It is also emphasized that generative tasks are not exclusive to decoder-only models, and text-to-text models might be an architecture that needs to be explored more and potentially promising. For example, instruction-tuned versions of T5 with 3B parameters (m20, m22) provide higher results than decoder-only models, almost three times larger, applying to both datasets.
When explicitly comparing precision and recall values, which can be critical depending on the application, the optimal choice of models may vary. Alpaca (m39) and LLaMA 2-CHAT-7B (m40) demonstrate proficiency in precision, while Flan-T5-XXL (m21) and T0++ (m23) excel in recall. Also, BioLORD (m07-m09) models, with 110 million parameters, exhibit precision and AUC score differences equal to or less than 3 points compared to models with a number of parameters between 7 billion and 11 billion.
With respect to the datasets, the MS-CXR dataset poses more challenges for the models than the transcriptions dataset. This difference in difficulty could be attributed to factors such as (i) the format of the data, (ii) the fact that the reports are from different medical subfields, and (iii) the length of the reports. Nevertheless, the highest performers per approach consistently show results indicative of clinical knowledge or clinical notions.
5.5 The groups of generative models: LLaMA and GPT

### v2

@@ 5 Results and Discussion
The main findings are outlined below. For comparability, AUC scores reported in this section correspond to evaluating the eleven class-reduced version of Transcriptions (see section 4.1.2). In addition, to ensure the robustness of the results, bootstrapping with iterations is applied to each experiment. Supplementary results are found in Appendix D.
5.1 Text classification analysis
| Dataset | Metric | CES | NLI | MCQA | ||||||
| Model | Score | Prompt | Pooling | Model | Score | Prompt | Model | Score | ||
| MS-CXR | Accuracy | BioLORD-STAMB2-v1-STS2 | x | Avg. | RoBERTa-MNLI | x | T0++ | |||
| F1-score | BioLORD-STAMB2-v1-STS2 | Avg. | RoBERTa-MNLI | x | T0++ | |||||
| Precision | BioLORD-PMB | CLS | RoBERTa-MNLI | x | Alpaca 7B | |||||
| Recall | BioLORD-STAMB2-v1-STS2 | Avg. | RoBERTa-MNLI | x | T0++ | |||||
@@ 5.1 Text classification analysis
| Transcriptions | AUC score | BioLORD-STAMB2-v1-STS2 | x | Avg. | BART Large-MNLI | x | Flan-T5-XXL | |||
The highest F1 and AUC scores are achieved with the largest instruction-tuned T5 models, i.e., Flan-T5 (m19-m21) and T0 (m22-m23), as shown in Table 2 and Fig. 2. Some of these scores are above in the F1-score and in the AUC score. These instruction-tuned T5 models in question stand among all models considered, ranking within the top 10 highest-performing models in both datasets. Nevertheless, the optimal choice of models may vary when considering precision as the target metric, where Alpaca (m39) and LLaMA 2-CHAT-7B (m40) demonstrate high competence.
Conversely, the lowest F1 and AUC scores are paradoxically obtained with the base T5 models (m14-m17), as evidenced in Fig. 2. These models, along with their clinically fine-tuned versions (m24-m25), rank within the top 10 lowest-performing models on both datasets. Moreover, the (1/1) and (6/8) of the models underperforming a random evaluator in Transcriptions and MS-CXR datasets, respectively, belong to base and clinically fine-tuned T5 models.
Delving into each approach, BioLORD models (m07-m09) are consistently the best choice for the contextual embedding similarity approach. MS-CXR dataset is relatively more complex than the Transcriptions dataset for these models, as reflected by their ranking in performance: 11th versus 3rd place, respectively. BART Large-MNLI (m13) represents the best overall model for the NLI approach. For both datasets, BART Large-MNLI is included in the top 10 highest-performing models, while RoBERTa-MNLI (m12) only does so for MS-CXR dataset. For the multiple choice QA approach, instruction-tuned models stand out, which include instruction-tuned LLaMa models (m39-m40) to the aforementioned instruction-tuned T5 models. Notably, LLaMA 2-CHAT-7B (m40) is within the top 10 highest-performing models in both datasets. These highest performers per approach consistently show results indicative of clinical knowledge or clinical notions.
@@
The results of instruction-tuned T5 models support the feasibility of representing discriminative tasks as generative ones by framing them as instructions. These results also underline that generative tasks are not exclusive to decoder-only models, and text-to-text models may be a promising architecture to explore further. For example, versions of T5 tuned to instructions with 3B parameters (m20, m22) provide superior results to decoder-only models, almost three times larger, on both evaluated datasets.
Model size – More parameters alone do not always translate into better results
The experiments yield findings questioning the claim that larger models consistently deliver superior performance. The performance of the models as a function of the logarithm of their size is depicted in Fig. 3. Testing for monotonic relationships via Spearman’s correlation is only reported for the multiple-choice QA approach due to the lack of diversity in sizes or number of samples. There is insufficient evidence to conclude that the Spearman’s correlation between size and performance is statistically significant in either dataset.
The trend of performance improvement with increasing size is almost nonexistent in the contextual embedding similarity approach. As seen in Fig. 2, for instance, models such as SapBERT (m06) and BioLORD models (m07-m09), which excel in this approach, outperform even three times larger models on both datasets. Within the same models, the deltas in performance associated with increasing the number of parameters are inconclusive. BERT (m01) marginally outperforms BERT (m00) on the Transcriptions dataset, whereas the opposite is observed in all metrics on the MS-CXR dataset. BiomedBERT-large (abstracts only) (m04) surpasses, albeit marginally, BioMedBERT (abstracts only) (m03) on both evaluated datasets, excluding in precision. Furthermore, performance gains are evidenced when more training data is used, as shown by comparing BioMedBERT (abstracts only) and BiomedBERT (abstracts + full text) (m02).
Similarly, the effect of increasing the size on performance is not sufficiently clear or strong in the multiple-choice question answering approach. While positive Spearman’s correlations are obtained, there is insufficient evidence to deem them statistically significant. None of the p-values are , so the null hypothesis that the two variables have no ordinal correlation cannot be rejected. Within T5 models, the effect is minimal or inconsistent when considering their non-instruction-tuned versions (m14-m17, m24-m25). Within the instruction-tuned T5 models (m18-m23), a consistent positive effect of size on performance is observed for both FlanT5 and T0 models on Transcriptions and only for the latter on MS-CXR.
On the other hand, the results in NLI align with the expectation that larger models lead to better performance. However, more models are needed to draw a (solid) conclusion. To have a notion about the lower bound in performance between NLI-DeBERTa (m11) and the largest models evaluated, the difference between the highest and the lowest values obtained among the evaluated prompts, respectively, is calculated. Considering all metrics, these differences range between on MS-CXR, having thus that the largest models always lead to performance improvement. Reaching the same conclusion on the Transcriptions dataset is not straightforward, given the results obtained for RoBERTa-MNLI (m12), as depicted in Fig. 2. This model’s performance is closer to the performance of NLI-DeBERTa than to that of BART Large-MNLI (m13).
Altogether, the results do not provide sufficient evidence that only increasing the model size, in number of parameters, leads to an improvement in performance, whether comparing different or the same models. Although model size may be a relevant factor in determining performance, it is hypothesized that training data and objectives are more decisive in small pre-trained language models. This hypothesis aligns with findings in [10] and [12]. Expanding the sample size and diversity could be essential to validate these observations, considering a minimum of 30 or 35 models per approach.
Model domain – More than a specialized domain; model architecture, training data, and training objective
Current medical datasets remain relatively small compared to those of the general domain, covering only a tiny region of the medical knowledge space [84]. Domain specialization of models using only one of these datasets in question may limit their generalization ability [45].
The effectiveness of domain specialization in improving performance is not evident in the contextual embedding similarity approach, as displayed in Fig. 2. The domain-specific models considered in this approach are Bio+Clinical BERT (m10), BiomedBERT models (m02-m04), and SciBERT (m05). Bio+Clinical BERT achieves lower scores than expected, positioning around the middle of the performance ranking for this approach. Similarly, some of the BiomedBERT models are outperformed by BERT (m00) and BERT (m01), their general domain counterparts. These findings, present in both datasets, challenge the superiority of domain-specific models over general domain ones in the task being evaluated via contextual embedding similarity.
Although existing, evidence supporting the effectiveness of domain specialization is still limited and unclear in the multiple-choice question answering approach. The models to be compared are T5 models (m14-m15) versus their clinical specialized versions (m24-m25), and Alpaca (m39) versus MedAlpaca (m52). Differences between ClinicalT5 and T5 models are and in AUC scores and and in F1-scores. Similarly, differences between MedAlpaca and Alpaca are in AUC scores and in F1 scores. Due to these values, it can not be clearly stated that domain specialization positively impacts performance.
Considering the insights discussed and the remarkable performance of BioLORD (m07-m09) models, SapBERT (m06), Flan-T5 (m18-m21) models, and T0 (m22-m23) models in their respective approaches, the training data, training objectives, and model architectures are possibly critical in determining model generalization. Continual pre-training for named entity recognition or medical entity linkage using contrastive learning on UMLS data is likely one of the factors for the success of SapBERT and BioLORD models. Likewise, employing instruction-tuned text-to-text models represents a compelling approach to achieving high performance in multiple-choice QA. Due to the impossibility of concluding on the NLI approach, expanding the analysis to incorporate domain-specialized NLI models in biomedical and clinical domains could be valuable.
Prompting and instruction-tuning key to model performance
One of the central points of the study is to analyze the influence of prompting on the models and text classification approaches under investigation. Prompting impact is quantified as the difference in performance resulting from the prompt usage, with positive values indicating an improvement, in contextual embedding similarity and NLI. In multiple-choice QA, this impact is calculated as the variation in performance, expressed in standard deviations, when different instructions are used. The resulting distributions are shown in Fig. 4.
Using a prompt does not always confer benefits in contextual embedding similarity, as reflected by Fig. 4. On Transcriptions, the average impact on the AUC score is points, with values ranging from to . Using any of the proposed prompts improves performance for of the model + pooling strategy combinations. In contrast, none of these prompts led to AUC score improvements for BioLORD-PMB, BiomedBERT models, BERT, and SciBERT. On MS-CXR, the impact of the prompt on performance is more positive on average, albeit with more variability. The average impact on the F1-score is points, with values ranging from to . Similar values are reported on accuracy, precision, and recall. Employing any of the proposed prompts represents benefits for the to of the model + pooling strategy combinations, depending on the metric. The performance of BioMedBERT (abstracts only) and BiomedBERT-large (abstracts only) is enhanced with any of the prompts, whereas the performance of the BioLORD models and Bio+Clinical BERT is hindered.
More consistent benefits are observed than in contextual embedding similarity when examining the prompt impact in the NLI approach. On Transcriptions, any of the proposed prompts yields performance improvements, profiting larger models the most from its usage. The average impact on the AUC score is for BART Large-MNLI, for NLI-DeBERTa, and for RoBERTa-MNLI. On MS-CXR, using a prompt only sometimes results in gains, particularly for NLI-DeBERTa. For this model, the average impact on the F1-score is ; while for BART Large-MNLI and RoBERTa-MNLI is and , respectively. Moreover, positive prompt impacts are only observed on precision for NLI-DeBERTa. In both datasets, there are certain prompts with a high positive impact, whereas others do not, mostly independent of the model.
Similarly, prompt importance is also evident in the multiple-choice question answering approach, given its significant observed influence on model performance. The proportion of models performing better than a random evaluator (AUC score ) on Transcriptions increases from to with appropriate prompts. Similarly, the proportion of better than a random evaluator (F1-score ) on MS-CXR rises from to . Prompting importance is thus highlighted not only by the high performance achieved but also by the brittleness of the models. The latter is reflected by the variability in Fig. 4, and further supported by Figs. 17 and 18 in Appendix D. Between datasets, the highest sensitivity to the prompt is found when evaluating Transcriptions, such that, with certain prompts, the instruction-tuned models yield similar results to their base counterparts. Overall, no single prompt works universally well for all models.
Regarding instruction-tuning, these models generally outperform their non instruction-tuned counterparts. The instruction-tuned T5 versions, whether T0 or Flan-T5, in any size considered, exhibit superior performance than their base counterparts. Instruction-tuning also improves performance consistently for the LLaMA models, whereas this is not always the case for other generative models: MPT and GPT-J are exceptions on the Transcriptions dataset and Falcon on the MS-CXR dataset. Overall, this tuning technique represents a gain, with an average increase of points in the AUC score and points in the F1-score.
Summarizing, the results endorse the crucial role of the prompt and its wording in the model’s performance, with both positive and negative effects presented. Consequently, we advocate using prompts and advanced prompting techniques to guide the model toward better results. This process should also not be limited to a single prompt due to the observed and well-known phenomenon of prompt brittleness [13]. Regarding instruction tuning, this technique proves to be beneficial for the models. More details on the prompt impact can be found in Figs. 20 and 19 in Appendix D.
5.2 Conditional text generation analysis
LLaMA models (m38-m39) stand out as the ones with the highest predictive capacity among the models evaluated. Particularly, LLaMA 2-7B (m38) is the highest performer, with a mean perplexity of when including the BOS token and when not. LLaMA models are also notable for the low standard deviation of their mean, with approximate values of and depending on the BOS token usage. These standard deviations indicate higher confidence in the estimated value of the mean.
Conversely, BioGPT models (m48-m49) are the models with the most significant difficulty in comprehending the dataset. BioGPT (m48), the lowest performer, presents a mean perplexity of when including the BOS token and when not. The variability on the mean of these models is among the highest observed, with approximate standard deviations of and depending on the BOS token usage. These results are paradoxical considering that BioGPT is domain-specific while LLaMA 2-7B is not.
Similarly to previous findings for text classification, domain specialization does not necessarily imply surpassing general domain models. For the medium-size domain-specific models, it is observed that BioGPT (m48) does not outperform any of the general domain models, while GPT-2-PubMed Medium (m46) does. For the large size models, domain specialization proves beneficial; whereas for the XL and XXL sizes, neither Galactica (m50-m51) nor BioGPT-Large (m49) clearly outperforms general domain models. Consequently, the only specialized models that prove advantageous are GPT-2-PubMed (m46-m47).
On the other hand, increasing the model size contributes to improved performances, regardless of whether or not the BOS token is included. A slight performance improvement is also observed for the second versions (m42, m44) versus the first versions (m41, m43) of OpenLLaMA. This improvement is on average and points on the perplexity for the 3B and 7B parameter versions, respectively. Considering that the difference between these versions of OpenLLaMA is the dataset used for pre-training, the results obtained for conditional text generation do not contradict those for text classification.
Furthermore, the standard deviations of perplexity reveal the presence of exceptionally challenging samples for the models, that is, outliers, which is visually depicted in Fig. 21 in Appendix D. Moderate outliers, above quantile by times the IQR, represent between and of the data, with BioGPT models having the highest percentages. Extreme outliers, above quantile by three times the IQR, make up between and of the data, with most models exhibiting percentages around and .
Groups of generative models – LLaMA and GPT-2